Reports 1-1 of 1 Clear search Modify search
fabian.arellano - 17:25 Tuesday 17 November 2020 (15474) Print this report
Payload inpection on 16-11-2020.

With Hiarata-san and Washimi-san.

Preparation from the control room

  • Using the IM picomotors I changed the IM-Y from around -3,000 urad to around zero.
  • Using the Fishing Rod I moved F0 keystone from around -1,900 um to around -1,600 um (nominal).
  • Using 20,000 counts of coil magnet actuator I set the F1 keystone a reasonable distance from -100 um (nominal). The FR is jammed.
  • Using the FR I set the BF to around +100 (nominal).
  • I measured transfer functions of the IM. As before, all of the are alright except IM-V one.
  • The difference is that now, the IM-V TF measurement fails even when the BF is at 100 um. This is a very bad situation.

In situ preparation

  • Hirata-san opened the vacuum chamber top. We covered the contact surface of the cap at the rim with the usual white tape.
  • We locked the IP and closed the chamber.
  • We opened the +X side door and also covered the contact surface at rim with white tape.

Work inside the chamber

  • Hirata-san and I went inside the chamber and locked the IP.
  • We did visual inspection of the payload and of the visible part of the BF. Everything looked fine.
  • Still, I wanted to check whether the problem is produced by a cable. So, we locked the IM and IRM to the security structure and, at the Flying Saucer above the IM, we unclamped the kapton cable of the IM picomotors. One of the two M4 screws was completely removed. This was the most likely offender.
  • We measured the BF transfer function and it revealed an unhealthy system, which means this kapton cable is not the problem. We have not clamped the cable back yet
  • The white OSEM cables also look fine. It does not seem worth unclamping them to check.
  • Tomorrow we aim to check the inside of the BF with the fibre optic camera.
Search Help